OpenStudy (anonymous):

The area of a triangle is 10 square inches and its height is 5 inches. Find the base

OpenStudy (jdoe0001):

area of a triangle = \(A = \cfrac{1}{2}\times base \times height\)

OpenStudy (jdoe0001):

so, you have the Area, and you have the height \(\large 10 = \cfrac{1}{2}\times base \times 5\) now just solve for "base" to get ... well, the base :)

OpenStudy (anonymous):

Base is 4?

OpenStudy (goldphenoix):

Well. Let see. 1/2 * 4 * 5 = 10. Therefore, the base is 4. You are correct!

OpenStudy (anonymous):

Thanks!